A solid disk rotates in the horizontal plane at an angular velocity of 5.00 × 10-2 rad/s with respect to an axis perpendicular to the disk at its center. The moment of inertia of the disk is 0.15 kg.m2. From above, sand is dropped straight down onto this rotating disk, so that a thin uniform ring of sand is formed at a distance of 0.40 m from the axis. The sand in the ring has a mass of 0.50 kg. After all the sand is in place, what is the angular velocity of the disk?
Answer #1

Here when sand is dropped on ring no exrternal torque is acting on the disc so angular momentum remain conserved

Initial angular momentum = final angular momentum

I1w1 = I2w2

I1 = initial moment of inertia(MI) = MI of disc = 0.15kg-m^2

angular velocity w1 = 5x10^-2 rad/s

I2 = (MI of disc + MI of Ring)

I2 = (0.15 + MR^2)

I2 = (0.15+(0.50*0.40*0.40))

I2 = 0.23 kg-m^2

Now I1*w1 = I2*w2

0.15*5x10^-2 = 0.23*w2

w2 = 3.26 x 10^-2 rad/s

so disc will rotate with angular velocity 3.26x10^-2rad/s